3. Dopamine 800 mg in 500 mL DSW is running at 30 mL/hr. The client weighs 150 ihs Calculate the dose the patient isreceiving in meg/ke/min. Plese answar oSng mentional AnalysiS 4. A client weighing 50kg is prescribed a Dobutrex solution of 250mg in 500mL DSW ordered to titrate between 2.5 and 5 mca/ka/min. Pleag ansuver 5o0mL Dsw erderg tirate bedcen 2 A andkm.eoz on a. Determine the volumetric pump flow rate setting for the upper and lower limit.
Answer #1

ANSWER

3. Dose of patient receiving in mcg/kg/min.

         Formula

             Dose of patient receiving = Drug concentration * flow rate / patient weight * 60 (min/hr)

                   Here

                       Determine drug concentration

                                800mg in 500 ml = 1.6 ml.

                               convert to mcg     = 1.6 ml * 1000

                                                           = 1600 mcg.

                      Patient weight = 150lbs

                                   1 lbs = 0.45 kg

                                            = 68kg

                      Flow rate = 30 ml /hr.

            As per formula

            Dose of patient receiving = 1600 * 30 / 68 * 60

                                                              = 48000 / 4080

                                                              = 11.8 mcg/kg/min

                                                              = 12 mcg/kg/min.

4. Volumetric pump flow rate setting for upper and lower limit

       Formula

           Flow rate = Dose of patient receiving * patient weight * 60 (ml/hr) / Drug concentration.

           Dose of patient receiving at upper limit = 5 mcg/kg/min.

           Dose of patient receiving at lower limit = 2.5 mcg/kg/min.

                                          Patient weight = 50 kg.

                                                          Drug concentration = 250mg /500 ml

                                                                                          = 0.5 ml

                                                                                          = 0.5 * 1000

                                                                                         = 500 mcg.

                   As per formula

                                      Lower limit flow rate = 2.5 * 50 *60 / 500

                                                                      = 7500 / 500

                                                                    = 15 ml/hr.

                                     Upper limit flow rate = 5 *50*60 / 500

                                                                          = 15000/500

                                                                          = 30 ml/hr.